Solve without using laplace 1. 3y – 8y – 3y=0 y(0)=10, y(0)=0
Answer #1

Please see attached solution

Please like itSolve without using complace 3y- 84-3y o 3dy - 8 do - 34 20 duz (302-8– 377 - 0 AE 3 msm-3=0 3 m (9-1)m-3=0 3m-qmtm-3=0 3my=getataen 10=0, e trofee 10=9 te dy a es tu x - 7/3 + + 36 eu Tu y case, eset3684 o ac, e ° +360° - 0 = + 36 2 -(+96C +4=10 -(+96=0 loc, la 1921 G+4=10 G = 10- 9 34 Y = geiße ly. geta tek like it plean